# A Framework for Objective-Driven Dynamical Stochastic Fields

### v1

@@ 3.1 Objective-Driven Dynamical Stochastic Field: Minimizing the Objective Value
The key question discussed in this subsection is seemingly a contradiction: how does an entity learn to minimize its objective value while following the same time-invariant transition rule? Intuitively, one might consider the time-invariant behavior of “minimization” itself as an inherent nature of the entity, with the outcomes of this minimization representing its learned behavior. For example, consider a program designed to minimize a certain criterion; the program itself remains fixed. This implies that the entity can simulate a minimizer through fixed dynamics, thereby exhibiting objective-driven behavior through time-invariant mechanisms.
Let us first formulate such objective-driven dynamics first through the two-entity view (Section 2.1). Consider a system composed of entity $x$ and $x^{\prime}$, i.e., ${\Omega}={\mathcal{A}}\times{\mathcal{B}}\times{\mathcal{M}}\times{\mathcal{N}}$. Focusing on entity $x$, it has an objective operator ${\bm{\Gamma}}$ and the corresponding objective value ${\bar{\gamma}}$.
Notation. Consider that entity $x$ is associated with a set of generators ${{\mathbf{M}}}\in{{\mathcal{P}}}_{x}$; entity $x^{\prime}$ is associated with generators ${{\mathbf{N}}}\in{{\mathcal{P}}}_{x^{\prime}}$. We can see that ${{\mathbf{M}}}\in{{\mathcal{P}}}_{x}$ and ${{\mathbf{N}}}\in{{\mathcal{P}}}_{x^{\prime}}$ together determine the system dynamics via ${{\mathbf{G}}}={{\mathbf{M}}}+{{\mathbf{N}}}$, and thus they determine the objective value ${\bar{\gamma}}={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$. We focus on entity $x$ with generator ${{\mathbf{M}}}$, and thus we denote ${\bar{\gamma}}({{\mathbf{M}}}):={\bar{\gamma}}={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$ as the objective value obtained by entity $x$ in this setting. Then, we make the following definition about what it means for a dynamics to be objective-driven.
@@
We may view entity $x^{\prime}$ as the environment and entity $x$ as an agent. The objective operator ${\bm{\Gamma}}$ then describes the objective signal the environment provides to the agent. Therefore, the objective-driven dynamics ${{\mathbf{M}}}^{\star}$ can be understood as the time-invariant dynamics that learn to adapt to the environment in such a way that they minimize the given objective value.
We prove the existence of such ${{\mathbf{M}}}^{\star}$ under some conditions, leaving the generic case for future study.
@@
Consider the two-entity formulation described above for an ergodic system. If the objective operator ${\bm{\Gamma}}\colon{\widetilde{\mathcal{H}}}\to{\mathbb{R}}_{+}$ is non-negative and $\forall{{{\mathbf{N}}}\in{{\mathcal{P}}}_{x^{\prime}}}:\min_{{{\mathbf{M}}}\in{{\mathcal{P}}}_{x}}{\bar{\gamma}}({{\mathbf{M}}})=0$ for a finite ${{\mathcal{P}}}_{x}$, then there exists generator ${{\mathbf{M}}}^{\star}$ that simulates the minimization, such that
@@
The notion of objective-driven dynamics can be extended to the field formulation; that is, each $x\in{\mathcal{X}}$ is associated with a local objective operator ${\bm{\Gamma}}_{x}$, and the local dynamics ${{\mathbf{G}}}(x)$ simulates the minimization of the local objective value ${\bar{\gamma}}(x)={\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$.
@@
A objective operator ${\bm{\Gamma}}_{x}=\gamma_{{\omega}^{\prime}{\omega}}{{\check{\bm{\eta}}}}^{{\omega}^{\prime}{\omega}}$ is local to $x\in{\mathcal{X}}$ if
@@
Consider a dynamical stochastic field where each entity $x\in{\mathcal{X}}$ is associated with a local objective operator ${\bm{\Gamma}}_{x}$ and a set of local infinitesimal generators ${{\mathcal{P}}}_{x}$. A objective-driven dynamical stochastic field is formed by objective-driven dynamics w.r.t. ${{\mathcal{P}}}_{x}$ that simulates the “minimization” of the local objective value ${\bar{\gamma}}_{x}={\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$.
@@
The immediate question is: What exactly is this gradient? Next, we address this question by giving the explicit formula for the gradient.
@@ 3.2 The Gradient Formula
Let us imagine the system suddenly varies the transition rate $G^{{\omega}^{\prime}}_{{\omega}}$ a bit. This change would affect the probability that the system jumps to configuration ${\omega}^{\prime}$ from configuration ${\omega}$. Therefore, it should be related to ${\hat{\bm{e}}}_{{\omega}^{\prime}{\omega}}\in{\widetilde{\mathcal{H}}}$ in some ways. It turns out to be useful to first define the following projection operator and its resulting subspace of ${\mathcal{H}}$.
@@
${\mathbf{\Pi}}\ {\hat{\bm{e}}}_{{\omega}^{\prime}{\omega}}={\hat{\bm{e}}}_{{\omega}^{\prime}}-{\hat{\bm{e}}}_{{\omega}}.$ (3.11)
@@
The subspace ${\Delta\mathcal{H}}$ is the span of vectors of the form ${\hat{\bm{e}}}_{{\omega}^{\prime}}-{\hat{\bm{e}}}_{{\omega}}$. Moreover, it admits an alternative representation, as shown in the following result.
@@
${\Delta\mathcal{H}}={\mathbf{\Pi}}\ {\widetilde{\mathcal{H}}}=\left\{{\hat{\bm{\varphi}}}\in{\mathcal{H}}\mid\sum_{{\omega}}{\varphi}^{\omega}=0\right\}.$ (3.13)
There is an important fact about the projection operator: it projects the lifted generator ${\widetilde{\mathbf{G}}}\colon{\mathcal{H}}\to{\widetilde{\mathcal{H}}}$ back to the original generator ${{\mathbf{G}}}$.
@@
Next, given that ${\Delta\mathcal{H}}$ involves the difference for a transition step at the current moment, it would be essential to examine the long-term behavior of such a step. The following operator plays a crucial role in capturing this long-term behavior.
@@
Figure 2: A commutative diagram illustrating the relationships among spaces ${\Delta\mathcal{H}},{\mathcal{H}},{\widetilde{\mathcal{H}}}$, and linear operators ${{\mathbf{G}}},{\widetilde{\mathbf{G}}},{\mathbf{\Pi}}$, and ${{\mathbf{S}}}$.
@@
There is one last step before we can calculate the gradient, that is we must first clarify the parameter with respect to which we are differentiating. Recall that we define the infinitesimal generators ${{\mathbf{G}}}$ by enumerating all possible behaviors of this linear operator, i.e., ${{\mathbf{G}}}={\hat{\bm{e}}}_{{\omega}^{\prime}}G^{{\omega}^{\prime}}_{{\omega}}{{\check{\bm{\eta}}}}^{{\omega}}$ (Definition 2.4). However, for it to be considered as a valid infinitesimal generator, it must satisfy a constraint $\sum_{{\omega}^{\prime}}G^{{\omega}^{\prime}}_{{\omega}}=0$.
@@
${{\mathbf{G}}}={\mathbf{\Pi}}\ {{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}G^{{\omega}^{\prime}}_{{\omega}},$ (3.16)
@@
Intuitively, the parameterization above can be understood as follows. Suppose we vary $G^{{\omega}^{\prime}}_{{\omega}}$ in ${{\mathbf{G}}}={\mathbf{\Pi}}\ {{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}G^{{\omega}^{\prime}}_{{\omega}}$. This variation would be equivalent to varying both $G^{{\omega}^{\prime}}_{{\omega}}$ and $G^{{\omega}}_{{\omega}}$ in ${{\mathbf{G}}}={\hat{\bm{e}}}_{{\omega}^{\prime}}G^{{\omega}^{\prime}}_{{\omega}}{{\check{\bm{\eta}}}}^{{\omega}}$, at the same time, such that ${{\mathbf{G}}}$ remains a valid generator. In other words, the parameterization above automatically satisfies the constraint $\sum_{{\omega}^{\prime}}G^{{\omega}^{\prime}}_{{\omega}}=0$.
With enough building blocks, we can finally calculate the gradient of the objective value ${\bar{\gamma}}={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$ w.r.t. the generator ${{\mathbf{G}}}$.
@@
Given an objective operator ${\bm{\Gamma}}\colon{\widetilde{\mathcal{H}}}\to{\mathbb{R}}$, assuming the system is ergodic, with the parameterization given by Definition 3.15 we have
$\frac{\partial{\bar{\gamma}}}{\partial G^{{\omega}^{\prime}}_{{\omega}}}={\bm{\Gamma}}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}{\hat{\bar{\bm{\varphi}}}}.$ (3.17)
@@
Notation. We denote ${{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu}\colon{\mathcal{H}}\to{\widetilde{\mathcal{H}}}$ as a linear operator that transforms ${\hat{\bm{e}}}_{\alpha\beta\mu\nu}$ to ${\hat{\bm{e}}}_{{\alpha^{\prime}\beta\mu^{\prime}\nu},{\alpha\beta\mu\nu}}$, regardless of $\nu$. Formally
${{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu}{\hat{\bm{e}}}_{\alpha^{\prime\prime}\beta^{\prime\prime}\mu^{\prime\prime}\nu^{\prime\prime}}=\begin{cases}{\hat{\bm{e}}}_{{\alpha^{\prime}\beta\mu^{\prime}\nu},{\alpha\beta\mu\nu}},\qquad&\text{if}\ \alpha\beta\mu=\alpha^{\prime\prime}\beta^{\prime\prime}\mu^{\prime\prime}\\ 0,&\text{otherwise}.\end{cases}$ (3.18)
Similarly we define ${{\widetilde{\mathbf{A}}}}_{\beta^{\prime}\nu^{\prime}}^{\alpha\beta\nu}$ for the other entity.
@@
${\widetilde{\mathbf{M}}}=M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}{{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu},\qquad\text{and}\qquad{\widetilde{\mathbf{N}}}=N^{\beta^{\prime}\nu^{\prime}}_{\alpha\beta\nu}{{\widetilde{\mathbf{A}}}}_{\beta^{\prime}\nu^{\prime}}^{\alpha\beta\nu}.$ (3.19)
@@
Given an objective operator ${\bm{\Gamma}}\colon{\widetilde{\mathcal{H}}}\to{\mathbb{R}}$, assuming the system is ergodic, with the parameterization of generator ${{\mathbf{M}}}$ given by Definition 3.17 we have
$\frac{\partial{\bar{\gamma}}}{\partial M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}}={\bm{\Gamma}}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu}{\hat{\bar{\bm{\varphi}}}}.$ (3.21)
Finally, we turn to the field formulation. It is easy to see that, given an entity $x\in{\mathcal{X}}$, we may view the rest of the field as the other entity, thus reducing the system to a two-entity view. Therefore, the action operator ${{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu}$ can be reformulated directly using the field notation. The important fact is that the action operator ${{\widetilde{\mathbf{A}}}}(x)_{{\omega}^{\prime}}^{{\omega}}$ changes the local configuration ${\omega}_{|{\Omega}_{x}}$ to ${\omega}^{\prime}_{|{\Omega}_{x}}$. Similar to previous discussions, given an arbitrary configuration ${\omega}^{\prime\prime}$, ${{\widetilde{\mathbf{A}}}}(x)_{{\omega}^{\prime}}^{{\omega}}$ would act on ${\hat{\bm{e}}}_{{\omega}^{\prime\prime}}$ only if ${\omega}^{\prime\prime}$ and ${\omega}$ agree on the local neighborhood ${{\mathcal{U}}}_{x}$. Then, the action would leave ${\Omega}/{\Omega}_{x}$ unchanged, resulting in ${\omega}^{\prime}_{|{\Omega}_{x}}\times{\omega}^{\prime\prime}_{|{\Omega}/{\Omega}_{x}}$. Note that we adopt the convention of automatic ordering: ${\omega}^{\prime\prime}_{|{\Omega}/{\Omega}_{x}}\times{\omega}^{\prime}_{|{\Omega}_{x}}$ refers to the same configuration as ${\omega}^{\prime}_{|{\Omega}_{x}}\times{\omega}^{\prime\prime}_{|{\Omega}/{\Omega}_{x}}$. This action operator is formally defined as follows.
Notation. We denote ${{\widetilde{\mathbf{A}}}}(x)_{{\omega}^{\prime}}^{{\omega}}\colon{\mathcal{H}}\to{\widetilde{\mathcal{H}}}$ as a linear operator that transforms ${\hat{\bm{e}}}_{{\omega}}$ to ${\hat{\bm{e}}}_{{{\omega}^{+}}{{\omega}}}$, where ${\omega}^{+}={\omega}^{\prime}_{|{\Omega}_{x}}\times{\omega}^{\prime\prime}_{|{\Omega}/{\Omega}_{x}}$, given that ${\omega}$ and ${\omega}^{\prime}$ agrees on ${\Omega}/{\Omega}_{x}$. Formally, it is defined as:
${{\widetilde{\mathbf{A}}}}(x)_{{\omega}^{\prime}}^{{\omega}}{\hat{\bm{e}}}_{{\omega}^{\prime\prime}}=\begin{cases}{\hat{\bm{e}}}_{{\omega}^{+}{\omega}},\quad&\text{if}\ {\omega}_{|{{\mathcal{U}}}_{x}}={\omega}^{\prime\prime}_{|{{\mathcal{U}}}_{x}}\ \text{and}\ \ {\omega}^{\prime}_{|{\Omega}/{\Omega}_{x}}={\omega}_{|{\Omega}/{\Omega}_{x}}\\ 0,&\text{otherwise}.\end{cases}$ (3.22)
@@
${{\mathbf{G}}}(x)={\mathbf{\Pi}}\ G(x)^{{\omega}^{\prime}}_{{\omega}}{{\widetilde{\mathbf{A}}}}(x)_{{\omega}^{\prime}}^{{\omega}}.$ (3.23)
@@
Given a local objective operator ${\bm{\Gamma}}_{x}\colon{\widetilde{\mathcal{H}}}\to{\mathbb{R}}$, assuming the system is ergodic, the gradient formula for ${\bar{\gamma}}(x)={\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$ w.r.t. local generator ${{\mathbf{G}}}(x)$ is
$\frac{\partial{\bar{\gamma}}(x)}{\partial G(x)_{{\omega}}^{{\omega}^{\prime}}}={\bm{\Gamma}}_{x}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){{\widetilde{\mathbf{A}}}}(x)^{{\omega}}_{{\omega}^{\prime}}{\hat{\bar{\bm{\varphi}}}}.$ (3.24)
The objective operators determine the behavior of the system, as each entity minimizes its objective value. However, if we want our field to perform some tasks, it is unclear what operators could give us the desired result, especially for such a complex, dynamical, stochastic, and decentralized system. This is of great importance and is our immediate focus.

### v2

@@ 3.1 Objective-Driven Dynamical Stochastic Field: Minimizing the Objective Value
The key question discussed in this subsection is seemingly contradictory: how does an entity learn to minimize its objective value while following the same time-invariant transition rule? Intuitively, one might consider the time-invariant behavior of “minimization” itself as an inherent nature of the entity, with the outcomes of this minimization representing its learned behavior. For example, consider a program designed to minimize a certain criterion; the program itself remains fixed. This implies that the entity can simulate a minimizer through fixed dynamics, thereby exhibiting objective-driven behavior through time-invariant mechanisms.
Let us first formulate such objective-driven dynamics first through the two-entity view (Section 2.1). Consider a system composed of entities $x$ and $x^{\prime}$, i.e., ${\Omega}={\mathcal{A}}\times{\mathcal{B}}\times{\mathcal{M}}\times{\mathcal{N}}$. Focusing on entity $x$, it has an objective operator $\bm{\Gamma}$} and the corresponding objective value $\bar{\gamma}$}.
Notation. Consider that entity $x$ is associated with a set of generators ${{\mathbf{M}}}\in{{\mathcal{P}}}_{x}$; entity $x^{\prime}$ is associated with generators ${{\mathbf{N}}}\in{{\mathcal{P}}}_{x^{\prime}}$. We can see that ${{\mathbf{M}}}\in{{\mathcal{P}}}_{x}$ and ${{\mathbf{N}}}\in{{\mathcal{P}}}_{x^{\prime}}$ together determine the system dynamics via ${{\mathbf{G}}}={{\mathbf{M}}}+{{\mathbf{N}}}$, and thus they determine the objective value ${\bar{\gamma}}={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$. We focus on entity $x$ with generator $\mathbf{M}$}}, and thus we denote ${\bar{\gamma}}({{\mathbf{M}}}):={\bar{\gamma}}={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$ as the objective value obtained by entity $x$ in this setting. Then, we make the following definition of what it means for dynamics to be objective-driven.
@@
We may view entity $x^{\prime}$ as the environment and entity $x$ as an agent. The objective operator $\bm{\Gamma}$} then describes the objective signal that the environment provides to the agent. Therefore, the objective-driven dynamics $\mathbf{M}$$^{\star}$ can be understood as the time-invariant dynamics that learn to adapt to the environment in such a way that they minimize the given objective value.
We prove the existence of such $\mathbf{M}$$^{\star}$ under certain conditions, leaving the generic case for future study.
@@
[…]
@@
The notion of objective-driven dynamics can be extended to the field formulation; that is, each $x\in{\mathcal{X}}$ is associated with a local objective operator ${\bm{\Gamma}}_{x}$, and the local dynamics ${{\mathbf{G}}}(x)$ simulate the minimization of the local objective value ${\bar{\gamma}}(x)={\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$.
@@
A objective operator ${\bm{\Gamma}}_{x}=\sum_{{\omega},{\omega}^{\prime}\in{\Omega}}\gamma_{{\omega}^{\prime}{\omega}}\langle{\omega}^{\prime}{\omega}\rvert$ is local to $x\in{\mathcal{X}}$ if
@@
Consider a dynamical stochastic field where each entity $x\in{\mathcal{X}}$ is associated with a local objective operator ${\bm{\Gamma}}_{x}$ and a set of local infinitesimal generators ${{\mathcal{P}}}_{x}$. A objective-driven dynamical stochastic field is formed by objective-driven dynamics w.r.t. ${{\mathcal{P}}}_{x}$ that simulates the “minimization” of the local objective value ${\bar{\gamma}}_{x}={\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$.
@@
The immediate question is: What exactly is this gradient? Next, we address this question by providing the explicit formula for the gradient.
@@ 3.2 The Gradient Formula
Let us imagine the system suddenly varies the transition rate $G^{{\omega}^{\prime}}_{{\omega}}$ slightly. This change would affect the probability that the system jumps to configuration $\omega$$^{\prime}$ from configuration $\omega$}. Therefore, it should be related to $\lvert{\omega}^{\prime}{\omega}\rangle\in{\widetilde{\mathcal{H}}}$ in some ways. It turns out to be useful to first define the following projection operator and its resulting subspace of $\mathcal{H}$}.
@@
${\mathbf{\Pi}}\ \lvert{\omega}^{\prime}{\omega}\rangle=\lvert{\omega}^{\prime}\rangle-\lvert{\omega}\rangle.$ (3.11)
@@
The subspace $\Delta\mathcal{H}$} is the span of vectors of the form $\lvert{\omega}^{\prime}\rangle-\lvert{\omega}\rangle$. Moreover, it admits an alternative representation, as shown in the following result.
@@
Write ${\lvert\varphi\rangle}=\sum_{{\omega}}{\varphi}^{\omega}\lvert{\omega}\rangle$, and we have
${\Delta\mathcal{H}}={\mathbf{\Pi}}\ {\widetilde{\mathcal{H}}}=\left\{{\lvert\varphi\rangle}\in{\mathcal{H}}\mid\sum_{{\omega}}{\varphi}^{\omega}=0\right\}.$ (3.13)
[…]
@@
Next, given that $\Delta\mathcal{H}$} involves the difference for a transition step at the current moment, it is essential to examine the long-term behavior of such a step. The following operator plays a crucial role in capturing this long-term behavior.
@@
Figure 2: A commutative diagram illustrating the relationships among spaces $\Delta\mathcal{H}$$\mathcal{H}$$\widetilde{\mathcal{H}}$}, and linear operators ${{\mathbf{G}}},{\widetilde{\mathbf{G}}},{\mathbf{\Pi}}$, and $\mathbf{S}$}}.
@@
There is one last step before we can calculate the gradient: we must first clarify the parameter with respect to which we are differentiating. Recall that we define the infinitesimal generators $\mathbf{G}$}} by enumerating all possible behaviors of this linear operator, i.e., ${{\mathbf{G}}}=\sum_{{\omega}{\omega}^{\prime}}\lvert{\omega}^{\prime}\rangle G^{{\omega}^{\prime}}_{{\omega}}\langle{\omega}\rvert$ (Definition 2.4). However, for it to be considered a valid infinitesimal generator, it must satisfy a constraint $\sum_{{\omega}^{\prime}}G^{{\omega}^{\prime}}_{{\omega}}=0$.
@@
${{\mathbf{G}}}=\sum_{{\omega},{\omega}^{\prime}\in{\Omega}}{\mathbf{\Pi}}\ {{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}G^{{\omega}^{\prime}}_{{\omega}},$ (3.16)
@@
Intuitively, the parameterization above can be understood as follows. Suppose we vary $G^{{\omega}^{\prime}}_{{\omega}}$ in ${{\mathbf{G}}}=\sum_{{\omega}{\omega}^{\prime}}{\mathbf{\Pi}}\ {{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}G^{{\omega}^{\prime}}_{{\omega}}$. This variation would be equivalent to varying both $G^{{\omega}^{\prime}}_{{\omega}}$ and $G^{{\omega}}_{{\omega}}$ in ${{\mathbf{G}}}=\sum_{{\omega}{\omega}^{\prime}}\lvert{\omega}^{\prime}\rangle G^{{\omega}^{\prime}}_{{\omega}}\langle{\omega}\rvert$, at the same time, such that $\mathbf{G}$}} remains a valid generator. In other words, the parameterization above automatically satisfies the constraint $\sum_{{\omega}^{\prime}}G^{{\omega}^{\prime}}_{{\omega}}=0$.
With enough building blocks, we can finally calculate the gradient of the objective value ${\bar{\gamma}}={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$ w.r.t. the generator $\mathbf{G}$}}.
@@
[…]
$\frac{\partial{\bar{\gamma}}}{\partial G^{{\omega}^{\prime}}_{{\omega}}}={\bm{\Gamma}}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){{\widetilde{\mathbf{A}}}}_{{\omega}^{\prime}}^{{\omega}}{\lvert\bar{\varphi}\rangle}.$ (3.17)
@@
Notation. We denote ${{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu}\colon{\mathcal{H}}\to{\widetilde{\mathcal{H}}}$ as a linear operator that transforms $\lvert\alpha\beta\mu\nu\rangle$ to $\lvert{\alpha^{\prime}\beta\mu^{\prime}\nu}\ {\alpha\beta\mu\nu}\rangle$, regardless of $\nu$. Formally
${{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu}\lvert\alpha^{\prime\prime}\beta^{\prime\prime}\mu^{\prime\prime}\nu\rangle=\begin{cases}\lvert{\alpha^{\prime}\beta\mu^{\prime}\nu}\ {\alpha\beta\mu\nu}\rangle,\qquad&\text{if}\ \alpha\beta\mu=\alpha^{\prime\prime}\beta^{\prime\prime}\mu^{\prime\prime}\\ 0,&\text{otherwise}.\end{cases}$ (3.18)
Similarly, we define ${{\widetilde{\mathbf{A}}}}_{\beta^{\prime}\nu^{\prime}}^{\alpha\beta\nu}$ for the other entity.
@@
${\widetilde{\mathbf{M}}}=\sum_{\alpha\beta\mu\alpha^{\prime}\mu^{\prime}}M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}{{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu},\qquad\text{and}\qquad{\widetilde{\mathbf{N}}}=\sum_{\alpha\beta\nu\alpha^{\prime}\nu^{\prime}}N^{\beta^{\prime}\nu^{\prime}}_{\alpha\beta\nu}{{\widetilde{\mathbf{A}}}}_{\beta^{\prime}\nu^{\prime}}^{\alpha\beta\nu}.$ (3.19)
@@
[…]
$\frac{\partial{\bar{\gamma}}}{\partial M^{\alpha^{\prime}\mu^{\prime}}_{\alpha\beta\mu}}={\bm{\Gamma}}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu}{\lvert\bar{\varphi}\rangle}.$ (3.21)
Finally, we turn to the field formulation. It is easy to see that, given an entity $x\in{\mathcal{X}}$, we may view the rest of the field as the other entity, thus reducing the system to a two-entity view. Therefore, the action operator ${{\widetilde{\mathbf{A}}}}_{\alpha^{\prime}\mu^{\prime}}^{\alpha\beta\mu}$ can be reformulated directly using the field notation. The important fact is that the action operator ${{\widetilde{\mathbf{A}}}}(x)_{{\omega}^{\prime}}^{{\omega}}$ changes the local configuration ${\omega}_{|{\Omega}_{x}}$ to ${\omega}^{\prime}_{|{\Omega}_{x}}$. Similar to previous discussions, given an arbitrary configuration $\omega$$^{\prime\prime}$, ${{\widetilde{\mathbf{A}}}}(x)_{{\omega}^{\prime}}^{{\omega}}$ would act on $\lvert{\omega}^{\prime\prime}\rangle$ only if $\omega$$^{\prime\prime}$ and $\omega$} agree on the local neighborhood ${{\mathcal{U}}}_{x}$. Then, the action would leave ${\Omega}/{\Omega}_{x}$ unchanged, resulting in ${\omega}^{\prime}_{|{\Omega}_{x}}\times{\omega}^{\prime\prime}_{|{\Omega}/{\Omega}_{x}}$. Note that we adopt the convention of automatic ordering: ${\omega}^{\prime\prime}_{|{\Omega}/{\Omega}_{x}}\times{\omega}^{\prime}_{|{\Omega}_{x}}$ refers to the same configuration as ${\omega}^{\prime}_{|{\Omega}_{x}}\times{\omega}^{\prime\prime}_{|{\Omega}/{\Omega}_{x}}$. This action operator is formally defined as follows.
Notation. We denote ${{\widetilde{\mathbf{A}}}}(x)_{{\omega}^{\prime}}^{{\omega}}\colon{\mathcal{H}}\to{\widetilde{\mathcal{H}}}$ as a linear operator that transforms $\lvert{\omega}\rangle$ to $\lvert{{\omega}^{+}}{{\omega}}\rangle$, where ${\omega}^{+}={\omega}^{\prime}_{|{\Omega}_{x}}\times{\omega}^{\prime\prime}_{|{\Omega}/{\Omega}_{x}}$, given that $\omega$} and $\omega$$^{\prime}$ agree on ${\Omega}/{\Omega}_{x}$. Formally, it is defined as:
${{\widetilde{\mathbf{A}}}}(x)_{{\omega}^{\prime}}^{{\omega}}\lvert{\omega}^{\prime\prime}\rangle=\begin{cases}\lvert{\omega}^{+}{\omega}\rangle,\quad&\text{if}\ {\omega}_{|{{\mathcal{U}}}_{x}}={\omega}^{\prime\prime}_{|{{\mathcal{U}}}_{x}}\ \text{and}\ \ {\omega}^{\prime}_{|{\Omega}/{\Omega}_{x}}={\omega}_{|{\Omega}/{\Omega}_{x}}\\ 0,&\text{otherwise}.\end{cases}$ (3.22)
@@
${{\mathbf{G}}}(x)=\sum_{{\omega},{\omega}^{\prime}\in{\Omega}}{\mathbf{\Pi}}\ G(x)^{{\omega}^{\prime}}_{{\omega}}{{\widetilde{\mathbf{A}}}}(x)_{{\omega}^{\prime}}^{{\omega}}.$ (3.23)
@@
Given a local objective operator ${\bm{\Gamma}}_{x}\colon{\widetilde{\mathcal{H}}}\to{\mathbb{R}}$, assuming the system is ergodic, the gradient formula for ${\bar{\gamma}}(x)={\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$ w.r.t. local generator ${{\mathbf{G}}}(x)$ is
$\frac{\partial{\bar{\gamma}}(x)}{\partial G(x)_{{\omega}}^{{\omega}^{\prime}}}={\bm{\Gamma}}_{x}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){{\widetilde{\mathbf{A}}}}(x)^{{\omega}}_{{\omega}^{\prime}}{\lvert\bar{\varphi}\rangle}.$ (3.24)
The objective operators determine the behavior of the system, as each entity minimizes its objective value. However, if we want our field to perform some tasks, it is unclear what operators could provide us with the desired result, especially for such a complex, dynamical, stochastic, and decentralized system. This is of great importance and is our immediate focus.